as a function of time is given by Fig. 2-36. 87. A


How do I solve this and what is the correct answer?
Answer #1

maximum decelration = - 5.8 m/s^2

Suppose She wants to stop:

u = 45 km/h = 45 x 1000m / 3600s = 12.5 m/s

a = - 5.8 m/s^2

t = 2s

distance needed to stop:

vf^2 - vi^2 = 2 a d

0^2 - 12.5^2 = 2(-5.8) d

d = 13.5 m

So it can stops within 28 m.

If she wants to cross:

for maximum acc. vf = 65 x 1000/3600 = 18.06 m/s

and vi = 12.5 m/s

18.06 = 12.5 + a(6)

a = 0.926 m/s^2


she had 2 sec.

lets see, how much distance she can cover.

d = ut + at^2 / 2

d = (12.5 x 2) + (0.926 x 2^2 /2 ) = 26.8 m

but if she want to cross within time then she have to cover 15+ 28 = 43 m.

which she can't.

so she must stop.